# Machine learning from limited data: Predicting biological dynamics under a time-varying external input

### v1

@@ Abstract
No electric field. We predict dynamics for $\mathbf{\hat{E}}=0$, which was not trained on, to assess the reservoir’s ability to interpolate unseen dynamics of its phase space on biological dynamics. The parallel architecture was utilized for this fully closed prediction. We use the same trained $\mathbf{W}_{\text{out}}$ but set $p=0$. Fig.4c,d demonstrates that the prediction was able to capture the mostly Brownian noise-like dynamics of the system, as evident in the spatiotemporal scales of the boundary fluctuations, without any macroscale polarization. Indeed, $H_{\text{prediction}}=0.56\pm 0.01$, which matches closely to that of the data ($H_{\text{data}}=0.55\pm 0.01$).
@@
Simulated transient. That RCs can predict the transients from sparse data motivates the question of whether the steady-states or parameter values had betrayed any information of the transients, or if the reservoir was able to interpolate the transients from the sparse transient observations alone. While studies have been conducted on the ability of RCs to anticipate transients in non-steady chaotic systems by providing training data of parameter values neighboring the transient regime, the mean lifetime of these transients obeyed a close relationship with the external parameter [17, 18]. Therefore, with ample training data of neighboring parameter values, the reservoir had information to assist in the interpolation or extrapolation to this unseen regime.
@@
We proceed to investigate the extent to which a reservoir can learn transients with sparse data by considering a simple linear system, which makes it easy to equip it with transient behavior that is independent of the steady-state. We consider a single time series $Q(t)$ that evolves as a charging or discharging capacitor in a damped circuit, equipped with a simple periodic forcing:
@@
The solutions to these equations (with $Q_{\text{charging}}(0)=0$ and $Q_{\text{discharging}}(0)=Q_{0}$) are then given by:
@@
We generate a training dataset with just the steady-states, and tag the discharged and charged states with $p=+1$ and $-1$, respectively (fig.4e). Here, $Q_{0}$ and $\tau$ were arbitrarily chosen to be 30 and 6, respectively. We let the reservoir predict the test data, which contains the transient. As seen in fig.4f, the reservoir can interpolate the timescale of the transient. Because this is a fully closed prediction, the reservoir has seen the transient only when it was synchronizing with the test data during the listening stage. This suggests that the reservoir has been able to learn not only the steady-states corresponding to the external parameter value, but also their transients, from sparse observations either in the training data or the test data, with a good choice of hyperparameters. In a dynamical systems framework, the reservoir is able to infer how to transition between coexisting basins of attractions in its phase space from very sparse observations.
.5 Conclusion

### v2

@@ III.3 Prediction of new steady states and transients
No electric field. We predict dynamics for $\mathbf{\hat{E}}=0$, which was not trained on, to assess the reservoir’s ability to interpolate unseen dynamics on biological dynamics. The parallel architecture was utilized for this fully closed prediction. We use the same trained $\mathbf{W}_{\text{out}}$ but set $p=0$. Fig.4c,d demonstrates that the prediction was able to capture the mostly Brownian noise-like dynamics of the system, as evident in the spatiotemporal scales of the boundary fluctuations, without any macroscale polarization. Indeed, $H_{\text{prediction}}=0.56\pm 0.01$, which matches closely to that of the data ($H_{\text{data}}=0.55\pm 0.01$).
@@
Simulated transient. That RCs can predict transients from sparse data motivates the question of whether the steady-states or parameter values had betrayed any information of the transients, or if the reservoir was able to interpolate the transients from the sparse transient observations alone. While studies have been conducted on the ability of RCs to anticipate transients in non-steady chaotic systems by providing training data of parameter values neighboring the transient regime, the mean lifetime of these transients obeyed a close relationship with the external parameter [18, 19]. Therefore, with ample training data of steady states corresponding to neighboring parameter values, the reservoir had information to assist in the interpolation or extrapolation to this unseen regime.
@@
We proceed to investigate the extent to which a reservoir can learn transients from sparse data by considering a simple linear system, which makes it easy to equip it with transients independent of steady-states. We consider a single time series $Q(t)$ that evolves as a charging or discharging capacitor in a damped circuit, equipped with a simple periodic forcing:
@@
The solutions to these equations (with $Q_{\text{charging}}(0)=0$ and $Q_{\text{discharging}}(0)=Q_{0}$) are given by:
@@
We generate a training dataset with just the steady-states, and tag the discharged and charged states with $p=+1$ and $-1$, respectively (fig.4e). $Q_{0}$ and $\tau$ were arbitrarily chosen to be 30 and 6, respectively. We let the reservoir predict the test data, which contains the transient. As seen in fig.4f, the reservoir can interpolate the timescale of the transient. Because this is a fully closed prediction, the reservoir has seen the transient only when it was synchronizing with the test data during the listening stage. This suggests that the reservoir has been able to learn not only the steady-states corresponding to the external parameter value, but also their transients, from sparse observations either in the training data or the test data, with a good choice of hyperparameters. In a dynamical systems framework, the reservoir is able to infer how to transition between coexisting basins of attractions in its phase space from very sparse observations.
IV Conclusion